We do a proof by induction to show the following inequality is true for all n∈N: 4n ≥ 4n

Let P(n) represent the statement: 4n ≥ 4n

For n=1: LHS = 41 = 4 and RHS = 4(1) = 4

Since 4≥4, P(1) is true.

Assume true for n=k: P(k)true means 4k ≥ 4k

Show true for n=k+1 assuming P(k):

Provide the correct proof.








1
Expert's answer
2021-05-07T14:13:37-0400

Given statement is-

"P(n):4^n\ge 4n ,\forall n\in N"


For "n=1, LHS= 4^1=4"

"RHS=4(1)=4=LHS"

Hence P(n) is true for n=1.


Let us assume that p(k) is true for some positive integer k.

"P(k):4^k\ge 4k~~~~~~-(1)"


We want to prove p(k+1) is true when p(k) is true-


"P(k+1): LHS= 4^{k+1}=4^k.4"

"RHS=4(k+1)=4k+4"


As "4^k\ge 4\Rightarrow 4^k.4\ge 4k+4"


So "P(k+1)" is true.


So given statement is true for all "n\in N."